### v1

@@ III-B Egress
Transactions are removed from the unconfirmed transaction pool for various reasons, such as inclusion in a block, limited transaction pool capacity, transaction expiry, fee priority or running out of gas (in Ethereum), replacement by a newer version that offers a higher fee, invalid or conflicting transaction or chain reorganization event at the node.
@@
We also set a parameter to remove older transactions in Neonpool, to make space for newer transactions. This is necessary because we do not perform any deletions on the bloom filter and the load in the bloom filter is bound to exceed the number of transactions it was initially dimensioned for. We consider two approaches: Firstly, the accumulated transactions can be periodically removed by clearing the filter. This period may be a fixed hourly interval or based on the number of transactions processed. Secondly, Decaying Bloom filters may be employed to randomly decay or age out transactions over time, to maintain an accurate representation of recent data while letting old data expire.
@@
Our work is a pioneering investigation into the feasibility and suitability of probabilistic data structures for transaction pool construction. Based on the core function of the transaction pool, we require a data structure with the following properties: can answer membership queries, can let old transactions expire to make way for more recent ones, and to ensure double spend protection key-value support is also required. While there are more than a dozen variants of bloom filters available in the literature, we start by using the most basic ones that meet our requirements and are widely understood. As proof-of-concept, we evaluate these in IV and gain greater insight into the theoretical limitations of each data structure. Further optimizations will be explored in an extension study.
III-C Scaling Neonpool
We next discuss a strategy to help Neonpool scale to increasing transaction loads: We initialize a counter to track total number of transactions in bloomTxFilter, and as soon as filter capacity is exceeded, additional bloom filters are generated recursively as per demand [31][32]. The user can tweak the filter capacity parameter as per the resources available at the node. The filters can then expire in the order of their age at defined intervals. New filters can be generated repeatedly if congestion persists.
IV Experiments, results and discussion
@@
IV-A Data set, implementation, and methodology
We record $ingress$ and $egress$ transactions in the transaction pool in JSON format (for raw transaction structure in Bitcoin and Ethereum, see [11] [33]) to allow us to reconstruct the transaction pool state at the client and replay network activity for simulation purposes. Our data set also includes all transactions (for Bitcoin inventory and Ethereum transaction message structure see [36] [34]) received over the network stored in CSV format. For Bitcoin, we run an instrumented version of Bitcoin Core modifying txmempool.cpp, to capture 10 million unique transactions (around 30 million transaction announcements over $\sim$30 days). Similarly, for Ethereum, we run an instrumented version of Geth, modifying txpool.go, to capture 10 million unique transactions (around 13 million transactions over $\sim$10 days).
We individually develop a simulation for Bitcoin mempool and Ethereum txpool using map data structures, with high-level pseudocode as described in Algo 2, and for Neonpool-BTC and Neonpool-ETH as described in Fig. 4 and 5. We replay transactions in each data set to reconstruct the Bitcoin mempool and Ethereum txpool over 30 and 10 days respectively. The simulated Bitcoin mempool and the Ethereum txpool act as the ground truth, and running it in parallel with Neonpool-BTC and Neonpool-ETH helps us evaluate how our scheme performs. Our dataset and C++ code are publicly accessible. For bloom filters and variants, we employ the Berkeley libbf library [20] which incorporates the $H3_{exp}$ hash functions.
We perform independent queries on Neonpool and the transaction pool at each $ingress$ transaction in our data set. The responses are recorded at each stage. The responses may diverge from the ground truth owing to the probabilistic nature of bloom filters. We discuss how these false positives affect our scheme and offer a quantitative analysis.
IV-B Evaluation metrics
@@
TABLE I: Neonpool-BTC performance for n=400k
@@
IV-C Error rates and memory utilization
IV-C1 Neonpool-BTC
The highest transaction volumes observed in the Bitcoin mempool to date is around 200k. We dimension bloomtxFilter to handle its double i.e. 400k transactions, because Neonpool delays the removal of transactions, as discussed below. Using Eq. 2 and Eq. 3, we dimension three filters of size 500 KB, 1 MB and 2 MB with 4M, 8M and 16M cells, having 7, 14, and 28 hash functions, and theoretical FPR of $038.2E-$ 03, $056.7E-$ 05, and $095.0E-$ 09 respectively.
We replay transaction events in the Bitcoin dataset. When Neonpoool-BTC is run without any transaction expiry mechanism, the transactions accumulate and quickly surpass the filter design capacity. Due to overloading in the filter, we get poor results. As shown in Tab. II, the 500 KB, 1 MB and 2 MB filters report a false positive rate (FPR) of $018.06E$ 01, $017.43E$ 01 and $016.82E$ 01, erroneously rejecting around 80.6%, 74.3% and 68.2% of transactions respectively. However, as no transactions are expired, the false negative rate (FNR) is zero.
We introduce expiry mechanisms to prevent overloading of the filters. We follow two approaches: 1. reset the bloom filter at fixed hourly intervals or once the count of transactions surpasses the number of transactions the filter was originally dimensioned for i.e. 400k in our case; 2. employ a decaying bloom filter that decrements a certain number of indices at random upon every insertion, hence mimicking expiry.
@@ IV-C1 Neonpool-BTC
Fig. 6 depicts in real-time the number of transactions stored in Neonpool-BTC with 1 MB filters and 24-hour expiry along with the FPR for 30 days (10 million unique transactions). We also plot the corresponding number of transactions in the Bitcoin mempool, the ground truth in our evaluation. The number of transactions closely tracks the pattern in the Bitcoin mempool, with an increasing offset, as egress transactions are retained in the filter until the expiry interval has lapsed.
@@
For instance, when the filter is cleared every 24 hours the average FPR, at $9.03-03$ is highest for the 500 kB filter, reducing to $031.94E-$ 03 and $046.22E-$ 04 as the filter size increases to 1 MB and further to 2 MB. For the 500 kB, 1 MB, and 2 MB filters, this translates to 286612 or $0.90\%$, 61577 or $0.19\%$ and 19735 or $0.06\%$ of transactions being erroneously rejected due to false positives. For each filter, there are around $0.18\%$ redundant transactions due to false negatives. Table I shows that as expected, as the expiry interval is reduced, the FPR improves, while the FNR deteriorates.
@@
Table I also shows that when the filter is cleared every 400k transactions, the average false positive rate at $039.80E-$ 03 is highest for the 500 kB filter and reducing to $031.7E-$ 03 and $036.00E-$ 03 as the filter size increases to 1 MB and further to 2 MB. For the 500 kB, 1 MB, and 2 MB filters. We observe $0.98\%$, $0.17\%$, and $0.06\%$ of transactions being erroneously rejected due to false positives. For each filter, there are around $0.15\%$ redundant transactions due to false negatives.
@@
Empirical false positive rates are significantly higher than the theoretical value, sometimes even more than an order of magnitude. We theorize the causes: first, multiple works have reported that false positive rates in real deployments are higher than theoretically computed [23] [22]. Researchers contend that this is because theoretical calculations assume that ”each hash transformation is perfect” [23] and that transactions ”are independent and uniformly distributed over all records” whereas real activity tends to be ”clumped” [22]. In this context, Bose et al. prove that Eq. 1 gives us a lower bound on the false positive rate [24].
@@
If we use a decaying bloom filter, we achieve vast improvements in terms of false positive rates. The average FPR at $051.19E-$ 05 is highest for the 500 kB filter and reduces to $068.53E-$ 06 and $061.98E-$ 06 as the filter size increases to 1 MB and further to 2 MB. For a decay factor of 128, the 500 kB, 1 MB, and 2 MB filters observe 498 or $0.0012\%$, 357 or $0.0009\%$ and 83 or $0.0002\%$ of transactions being erroneously rejected due to false positives and there are around $0.61\%$ redundant transactions due to false negatives. Table I shows that by increasing the decay factor the FPR and hence the number of erroneously rejected transactions reduce. This is because the decay average meets the insertion average, and the filter reaches a stable state. However on the flip side the false negative rate increases.
@@
Figure 7: Memory Usage of Bitcoin and Neonpool-BTC
IV-C2 Neonpool-ETH
@@
TABLE II: Neonpool-ETH: Performance for n=700,000
@@
We replay transaction events in the Ethereum dataset. First, we run Neonpoool-ETH without any transaction expiry mechanism. Thus transactions accumulate and quickly surpass the filter design capacity. Due to overloading in the filter, we get poor results. As shown in Table  II, the 500 KB, 1 MB and 2 MB filters report a FPR of $017.30E$ 01, $016.70E$ 01 and $016.07E$ 01, erroneously rejecting around 73.0%, 67.0% and 60.07% of transactions respectively. However, as no transactions are expired, the FNR is zero.
We introduce expiry mechanisms to prevent overloading of the filters. We follow two approaches: 1. reset the bloom filter at fixed hourly intervals or once the count of transactions surpasses the number of transactions the filter was originally dimensioned for i.e. 700k in our case; 2. employ a decaying bloom filter that decrements a certain number of indices at random upon every insertion, hence mimicking expiry.
Figure 8 depicts the number of transactions in Neonpool-ETH with 1 MB filters and 24-hour expiry along with the FPR for almost 10 days (10 million unique transactions). We also plot the corresponding number of transactions in the Ethereum txpool, the ground truth in our evaluation. The number of transactions closely tracks the pattern in the Ethereum txpool, with an increasing offset, as egress transactions are retained in the filter until the expiry interval has lapsed.
For instance, when the filter is cleared every 12 hours the average FPR, at $037.43E-$ 03 is highest for the 500 kB filter, reducing to $042.39E-$ 04 and $055.38E-$ 05 as the filter size increases to 1 MB and further to 2 MB. For the 500 kB, 1 MB, and 2 MB filters, this translates to 80758 or $0.74\%$, 2593 or $0.02\%$ and 584 or $0.0569\%$ of transactions being erroneously rejected, respectively. For each filter, there are over $1\%$ redundant transactions due to false negatives. Table II shows that as expected, as the expiry interval is reduced, the FPR improves, while the FNR deteriorates.
Table II also shows that when the filter is cleared every 700k transactions, the average FPR at $021.63E-$ 02 is highest for the 500 kB filter and reducing to $044.61E-$ 04 and $041.66E-$ 04 as the filter size increases to 1 MB and further to 2 MB. For the 500 kB, 1 MB, and 2 MB filters, this translates to $1.63\%$, $0.05\%$ and $0.017\%$ of transactions being erroneously rejected due to false positives, respectively. For each filter, there are around $0.64\%$ redundant transactions.
@@
If we use a decaying bloom filter, we achieve vast improvements in terms of false positive rates. The average FPR at $063.41E-$ 06 is highest for the 500 kB filter and reduces to $062.02E-$ 06 and $062.21E-$ 06 as the filter size increases to 1 MB and further to 2 MB. For a decay factor of 128, the 500 kB, 1 MB, and 2 MB filters observe 37 or $0.0003\%$, 22 or $0.0002\%$ and 24 or $0.0002\%$ of transactions being erroneously rejected due to false positives, respectively. For each filter, there are over $0.82\%$ redundant transactions due to false negatives. Table II shows that by increasing the decay factor the FPR and hence the number of erroneously rejected transactions reduce. This is because the decay average meets the insertion average, and the filter reaches a stable state. However on the flip side the false negative rate increases.
@@
Assuming each transaction is from a unique account, dstxFilter will have the same dimensions as bloomtxFilter and consequently similar FPR. Thus, for a 1 MB bloomtxFilter, rejecting around 0.0005% of valid transactions, the corresponding dstxFilter will also reject around 0.0005% of valid transactions. Neonpool-ETH achieves 99.999% fidelity, handling 400 MB of transactions in just 2 MB, as shown in Fig. 9.
Figure 9: Memory Usage of Ethereum and Neonpool-ETH
IV-D Computation time
We undertake some experiments to estimate the computation overhead of Neonpool. The map based transaction pool in both Bitcoin and Ethereum performs query, insertion, and deletion operations in $O(\log n)$ time, where $n$ represents the number of stored transactions. In Neonpool, bloom filters operate in constant time, $O(k)$, where $k$ is the number of hash functions.
We perform simulations on Raspberry Pi 4 - Broadcom BCM2711, Quad-core Cortex-A72 64-bit @ 1.8GHz with 8 GB RAM, and Jetson Nano Quad-core ARM Cortex-A57 64-bit @1.43 GHz MPCore processor 4 GB RAM. Table III shows the computation time in microseconds ($\mu$$s$), averaged over 1E06 iterations, for querying and inserting transactions.
Bitcoin, for querying and inserting a transaction and its inputs takes 20.2 $\mu$$s$ ($2\times(1.8+8.3)$ $\mu$$s$) on a Raspberry Pi 4 and 27.2 $\mu$$s$ on a Jetson Nano on average. Neonpool-BTC with bloomtxFilter and dstxFilter dimensioned at 1 MB with k=14 hash functions, cumulatively for query and insert will take 20.8 $\mu$$s$ ($2\times(4.4+6.0)$ $\mu$$s$) on a Raspberry Pi 4 and 26 $\mu$$s$ on a Jetson Nano on average to process one transaction.
Ethereum, for querying and inserting a transaction and the state information takes 17.4 $\mu$$s$ ($2\times(6.9+1.8)$ $\mu$$s$) on a Raspberry Pi 4 and 20.6 $\mu$$s$ on a Jetson Nano on average. Similarly, Neonpool-ETH with bloomtxFilter and dstxFilter, dimensioned at 1 MB and k=8 hash functions, cumulatively for query and insert, will take 12.4 $\mu$$s$ ($2\times(2.9+3.3)$ $\mu$$s$) on a Raspberry Pi 4 and 14.4 $\mu$$s$ on a Jetson Nano, on average.

### v2

@@ 3.2 Egress
In Bitcoin core and Ethereum, transactions are removed from the unconfirmed transaction pool for various reasons, such as inclusion in a block, limited transaction pool capacity, transaction expiry, fee priority, running out of gas (in Ethereum), replacement by a newer version that offers a higher fee, invalid or conflicting transaction, or chain reorganization event at the node.
In Bitcoin, transactions are automatically removed from the mempool after a default expiration period of 14 days, although this limit can be configured by the node operator [23]. In Ethereum, the transaction pool’s memory usage is configurable, and transactions expire when the number of pending transactions exceeds the default limit of 4,096 or when they remain unprocessed for more than 3 hours [24].
@@
Neonpool mimics these expiry mechanisms using Bloom filters, offering nodes configurable parameters for transaction expiry through periodic clearing and decaying Bloom Filters. This is necessary because we do not perform any deletions on the bloom filter, and the load in the bloom filter is bound to exceed the number of transactions it was initially dimensioned for. We consider two approaches: Firstly, the accumulated transactions can be periodically removed by clearing the filter. This period may be a fixed hourly interval or based on the number of transactions processed. Secondly, Decaying Bloom filters may be employed to randomly decay or age out transactions, to maintain an accurate representation of recent data while letting old data expire.
@@
Our work is a pioneering investigation into the feasibility and suitability of probabilistic data structures for transaction pool construction. Based on the core function of the transaction pool, we require a data structure with the following properties: it can answer membership queries, let old transactions expire to make way for more recent ones and ensure double spending protection. Key-value support is also required. While there are more than a dozen variants of bloom filters available in the literature, we start by using the most basic ones that meet our requirements and are widely understood. As a proof-of-concept, we evaluate these in § 4 and gain greater insight into the theoretical limitations of each data structure. Further optimizations will be explored in an extension study.
3.3 Scaling Neonpool
We present a strategy to enable Neonpool to handle increasing transaction loads effectively: To mitigate the risk of false positives during heavy network congestion, the system initializes a counter to track the total number of transactions inserted into the Bloom filter. When the transaction count exceeds the capacity of 200,000, additional auxiliary Bloom filters of the same size and capacity are dynamically instantiated, as per demand and expired in order of age [39][40]. This prevents overloading the filter, thereby maintaining acceptable false positive rates.
The proposed approach of recursively generating Bloom filters may result in increased computation time and memory overhead on lightweight devices. In cryptocurrencies like Bitcoin and Ethereum, there is an established practice of limiting transaction pool sizes and managing overflow by rejecting or expiring excess transactions. For instance, Bitcoin’s default mempool size is capped at 300 MB, while Ethereum employs a default limit of 4096 transactions in the transaction pool, with surplus transactions being evicted [47] [24]. Furthermore, users can customize transaction pool policies or disable transaction pools entirely to accommodate low-memory or low-computation environments.
Similarly, Neonpool allows nodes to configure and adjust memory and computational thresholds based on their specific capacity. If the memory or computational threshold is reached, older Bloom filters automatically expire as necessary. For example, if a node can efficiently manage only two concurrent filters without exceeding its computational resources, it can set this limit to ensure optimal performance.
The objective is to empower full-node users by optimizing resource utilization, providing enhanced flexibility, and granting greater control over resource allocation. This approach incentivizes altruistic participation of the full nodes, enabling them to actively participate in transaction verification and forwarding, thus contributing to the decentralization, security, and robustness of the ecosystem.
The scaling function of Neonpool is governed by Eq. 2 and is determined by the configuration of the Bloom filter, specifically the size of the bit array and the number of hash functions: False-Positive Rate (p): Eq. 2 shows that the bit array size (m) grows logarithmically with p. For example, reducing the false-positive probability from $10^{-3}$ to $10^{-6}$ increases the bit array size modestly. This logarithmic relationship allows for efficient use of resources while lowering false positives. Number of Transactions (n): Assuming the false-positive rate (p) is kept constant, the bit array size grows linearly with the number of transactions. As the number of transactions (n) increases, the size of the Bloom filter increases proportionally. If the number of transactions doubles, the Bloom filter size doubles, making it scalable as transaction pools grow.
Thus, Neonpool scales logarithmically with the desired false-positive probability and linearly with the number of transactions it tracks, ensuring efficient transaction pool management even in resource-constrained environments.
4 Experiments, results and discussion
@@
4.1 Data set, implementation, and methodology
We record $ingress$ and $egress$ transactions in the transaction pool in JSON format for Bitcoin and CSV for Ethereum (for raw transaction structure in Bitcoin and Ethereum, see [10] [41]) to allow us to reconstruct the transaction pool state at the client and replay network activity for simulation purposes. Our data set also includes all transactions (for Bitcoin inventory and Ethereum transaction message structure see [45] [42]) received over the network stored in CSV format. For Bitcoin, we run an instrumented version of Bitcoin Core modifying txmempool.cpp to capture 10 million unique transactions (around 30 million transaction announcements over $\sim$30 days). Similarly, for Ethereum, we run an instrumented version of Geth, modifying txpool.go, to capture 10 million unique transactions (around 13 million transactions over $\sim$10 days).
We develop simulations for Bitcoin mempool and Ethereum txpool using map data structures, with high-level pseudocode described in Algo 2, and for Neonpool-BTC and Neonpool-ETH as shown in Fig. 4(a) and 5(a). We replay transactions in each dataset to reconstruct the Bitcoin mempool and Ethereum txpool over 30 and 10 days, respectively. The simulated Bitcoin mempool and Ethereum txpool serve as the ground truth, running in parallel with Neonpool-BTC and Neonpool-ETH to evaluate our scheme’s performance.
Neonpool-BTC was implemented using C++, as Bitcoin Core, the most widely used Bitcoin client, is written in C++. Furthermore, C++ offers a robust standard library optimized for high performance, efficient memory management, and handling data-intensive operations with minimal overhead. For Bloom filters and variants, we use the comprehensive Berkeley libbf library [25] written in C++ and incorporate the $H3_{exp}$ hash functions.
For Neonpool-ETH, we reuse the code components developed for Neonpool-BTC. The language-agnostic nature of our approach ensures that Neonpool’s advantages are preserved even when ported to Go for integration with Ethereum. Our dataset and code are publicly accessible. [43]
We perform independent queries on Neonpool and the transaction pool at each $ingress$ transaction in our data set, and the responses are recorded. The responses may diverge from the ground truth owing to the probabilistic nature of bloom filters. We discuss how these false positives and negatives affect our scheme and offer a quantitative analysis.
4.2 Evaluation metrics
@@
Table 1: Neonpool-BTC performance for n=400k
@@
4.3 Error rates and memory utilization
4.3.1 Neonpool-BTC
The highest transaction volumes observed in the Bitcoin mempool to date are around 200k. We dimension bloomtxFilter to handle double that, i.e., 400k transactions, because Neonpool delays removing transactions, as discussed below. Using Eq. 2 and Eq. 3, we dimension filters of size 500 KB, 1 MB, and 2 MB with 4M, 8M, and 16M cells, having 7, 14, and 28 hash functions, and theoretical FPR of $038.2E-$ 03, $056.7E-$ 05, and $095.0E-$ 09 respectively.
We replay transaction events in the Bitcoin dataset. When Neonpool-BTC runs without a transaction expiry mechanism, transactions accumulate and quickly surpass the filter design capacity. Due to filter overloading, we get poor results. As shown in Tab. 2, the 500 KB, 1 MB, and 2 MB filters report a false positive rate (FPR) of $018.06E$ 01, $017.43E$ 01, and $016.82E$ 01, erroneously rejecting 80.6%, 74.3%, and 68.2% of transactions respectively. However, as no transactions expire, the false negative rate (FNR) is zero.
We introduce expiry mechanisms to prevent the filters from overloading. We follow two approaches: 1. reset the bloom filter at fixed hourly intervals or once the count of transactions surpasses the number of transactions the filter was originally dimensioned for i.e. 400k in our case; 2. employ a decaying bloom filter that decrements a certain number of indices at random upon every insertion, hence mimicking expiry.
@@ 4.3.1 Neonpool-BTC
Fig. 6 depicts in real-time the number of transactions stored in Neonpool-BTC with 1 MB filters and 24-hour expiry alongside the FPR for 30 days (10 million transactions). We also plot the corresponding number of transactions in the Bitcoin mempool, the ground truth in our evaluation. The number of transactions closely tracks the pattern in the Bitcoin mempool, with an increasing offset, as the filter retains egress transactions until the expiry interval lapses.
@@
For instance, when the filter is cleared every 24 hours the average FPR, at $9.03-03$ is highest for the 500 kB filter, reducing to $031.94E-$ 03 and $046.22E-$ 04 as the filter size increases to 1 MB and further to 2 MB. For the 500 kB, 1 MB, and 2 MB filters, this translates to 286612 or $0.90\%$, 61577 or $0.19\%$ and 19735 or $0.06\%$ of transactions being erroneously rejected due to false positives. For each filter, there are around $0.18\%$ redundant transactions due to false negatives. Tab. 1 shows that as the expiry interval is reduced, the FPR improves, while the FNR deteriorates.
@@
Tab. 1 shows that when the filter is cleared every 400k transactions, the average false positive rate at $039.80E-$ 03 is highest for the 500 kB filter and reducing to $031.7E-$ 03 and $036.00E-$ 03 as the filter size increases to 1 MB and further to 2 MB. For the 500 kB, 1 MB, and 2 MB filters. We observe $0.98\%$, $0.17\%$, and $0.06\%$ of transactions being erroneously rejected due to false positives. For each filter, there are around $0.15\%$ redundant transactions due to false negatives.
@@
Empirical false positive rates are significantly higher than the theoretical value, sometimes even more than an order of magnitude. We theorize the causes: first, multiple works have reported that false positive rates in real deployments are higher than theoretically computed [31] [30]. Researchers contend that this is because theoretical calculations assume that "each hash transformation is perfect" [31] and that transactions "are independent and uniformly distributed over all records" whereas real activity tends to be "clumped" [30]. In this context, Bose et al. prove that Eq. 1 gives us a lower bound on the false positive rate [32].
Hence, these deviations reflect practical realities rather than flaws in the theoretical framework. To address this, an effective strategy is to over-dimension the Bloom filter. This compensates for empirical deviations by ensuring false positive rates remain within acceptable bounds for specific applications. Such an approach aligns with established deployment practices, where configurations are optimized based on observed performance metrics rather than idealized theoretical models.
@@
If we use a decaying bloom filter, we achieve vast improvements in terms of false positive rates. The average FPR at $051.19E-$ 05 is highest for the 500 kB filter and reduces to $068.53E-$ 06 and $061.98E-$ 06 as the filter size increases to 1 MB and further to 2 MB. For a decay factor of 128, the 500 kB, 1 MB, and 2 MB filters observe 498 or $0.0012\%$, 357 or $0.0009\%$ and 83 or $0.0002\%$ of transactions being erroneously rejected due to false positives and there are around $0.61\%$ redundant transactions due to false negatives. Tab. 1 shows that by increasing the decay factor, the FPR and, hence, the number of erroneously rejected transactions reduce. This is because the decay average meets the insertion average, and the filter reaches a stable state. However, on the flip side, the false negative rate increases.
@@
Figure 7: Memory usage of mempool and Neonpool-BTC
4.3.2 Neonpool-ETH
@@
Table 2: Neonpool-ETH: Performance for n=700,000
@@
We replay transaction events in the Ethereum dataset. First, we run Neonpoool-ETH without any transaction expiry mechanism. Thus transactions accumulate and quickly surpass the filter design capacity. Due to overloading in the filter, we get poor results. As shown in Tab. 2, the 500 KB, 1 MB and 2 MB filters report a FPR of $017.30E$ 01, $016.70E$ 01 and $016.07E$ 01, erroneously rejecting around 73.0%, 67.0% and 60.07% of transactions respectively. However, as no transactions are expired, the FNR is zero.
We introduce expiry mechanisms to prevent the filters from overloading. We follow two approaches: 1. reset the bloom filter at fixed hourly intervals or once the count of transactions surpasses the number of transactions the filter was originally dimensioned for i.e. 700k in our case; 2. employ a decaying bloom filter that decrements a certain number of indices at random upon every insertion, hence mimicking expiry.
Fig. 8 depicts the number of transactions in Neonpool-ETH with 1 MB filters and 24-hour expiry along with the FPR for almost 10 days (10 million unique transactions). We also plot the corresponding number of transactions in the Ethereum txpool, the ground truth in our evaluation. The number of transactions closely tracks the pattern in the Ethereum txpool, with an increasing offset, as the filter retains egress transactions until the expiry interval lapses.
For instance, when the filter is cleared every 12 hours, the average FPR, at $037.43E-$ 03, is highest for the 500 kB filter, reducing to $042.39E-$ 04 and $055.38E-$ 05 as the filter size increases to 1 MB and further to 2 MB. For the 500 kB, 1 MB, and 2 MB filters, this translates to 80758 or $0.74\%$, 2593 or $0.02\%$ and 584 or $0.0569\%$ of transactions being erroneously rejected, respectively. For each filter, there are over $1\%$ redundant transactions due to false negatives. Tab. 2 shows that, as expected, as the expiry interval is reduced, the FPR improves while the FNR deteriorates.
Tab. 2 also shows that when the filter is cleared every 700k transactions, the average FPR at $021.63E-$ 02 is highest for the 500 kB filter and reducing to $044.61E-$ 04 and $041.66E-$ 04 as the filter size increases to 1 MB and further to 2 MB. For the 500 kB, 1 MB, and 2 MB filters, this translates to $1.63\%$, $0.05\%$ and $0.017\%$ of transactions being erroneously rejected due to false positives, respectively. For each filter, there are around $0.64\%$ redundant transactions.
@@
If we use a decaying bloom filter, we achieve vast improvements in terms of false positive rates. The average FPR at $063.41E-$ 06 is highest for the 500 kB filter and reduces to $062.02E-$ 06 and $062.21E-$ 06 as the filter size increases to 1 MB and further to 2 MB. For a decay factor of 128, the 500 kB, 1 MB, and 2 MB filters observe 37 or $0.0003\%$, 22 or $0.0002\%$ and 24 or $0.0002\%$ of transactions being erroneously rejected due to false positives, respectively. For each filter, there are over $0.82\%$ redundant transactions due to false negatives. Tab. 2 shows that by increasing the decay factor the FPR and hence the number of erroneously rejected transactions reduce. This is because the decay average meets the insertion average, and the filter reaches a stable state. However, on the flip side, the false negative rate increases.
@@
Assuming each transaction is from a unique account, dstxFilter will have the exact dimensions as bloomtxFilter and consequently similar FPR. Thus, for a 1 MB bloomtxFilter, rejecting around 0.0005% of valid transactions, the corresponding dstxFilter will also reject around 0.0005% of valid transactions. Neonpool-ETH achieves 99.999% fidelity, handling 400 MB of transactions in just 2 MB, as shown in Fig. 9.
Figure 9: Memory usage of txpool and Neonpool-ETH
4.4 Computation time
We conduct experiments to estimate the computation overhead of Neonpool. The map-based transaction pool in Bitcoin and Ethereum performs query, insertion, and deletion operations in $O(\log n)$ time, where $n$ is the number of stored transactions. In Neonpool, bloom filters operate in constant time, $O(k)$, where $k$ is the number of hash functions.
We perform simulations on Raspberry Pi 4 - Broadcom BCM2711, Quad-core Cortex-A72 64-bit @ 1.8GHz with 8 GB RAM, and Jetson Nano Quad-core ARM Cortex-A57 64-bit @1.43 GHz MPCore processor 4 GB RAM. Tab. 3 shows the computation time in microseconds ($\mu$$s$), averaged over 1E06 iterations, for querying and inserting transactions.
Neonpool is designed to avoid increasing computational demands, which directly translates to energy consumption. In fact, Neonpool reduces computation time compared to traditional transaction pools:
In Bitcoin, processing an incoming transaction begins with querying the transaction’s hash, followed by querying its inputs [^2] For simplicity, we assume each transaction has a single input. As the number of inputs increases, the number of queries increases proportionally in both Bitcoin Core and Neonpool-BTC.. If the transaction is new and has valid inputs, it is then added to the map along with its inputs. This process involves two queries and two insert operations. Similarly, in Neonpool-BTC, the transaction’s hash is queried first, followed by a query for its inputs. If the transaction is new and has valid inputs, it is added to Neonpool-BTC along with its inputs. This also involves two queries and two insert operations. The total computational cost for this process can be expressed as: $2\times(query\>time+insert\>time)$.
Bitcoin, for querying and inserting a single transaction and its inputs, takes 20.2 $\mu$$s$ ($2\times(1.8+8.3)$ $\mu$$s$) on a Raspberry Pi 4 and 27.2 $\mu$$s$ on a Jetson Nano on average. Neonpool-BTC with bloomtxFilter and dstxFilter dimensioned at 1 MB with k=14 hash functions, cumulatively for query and insert, takes 20.8 $\mu$$s$ ($2\times(4.4+6.0)$ $\mu$$s$) on a Raspberry Pi 4 and 26 $\mu$$s$ on a Jetson Nano, on average.
Similarly, in Ethereum, processing starts with querying the transaction’s hash, followed by querying the nonce if the transaction is fresh. If the transaction is new and has a valid nonce, it is added to the txpool, and the nonce is updated. This process also requires two queries and two insert operations. The computational cost for this process is the same: $2\times(query\>time+insert\>time)$.
Ethereum, for querying and inserting a single transaction and the state information, takes 17.4 $\mu$$s$ ($2\times(6.9+1.8)$ $\mu$$s$) on a Raspberry Pi 4 and 20.6 $\mu$$s$ on a Jetson Nano on average. Similarly, Neonpool-ETH with bloomtxFilter and dstxFilter, dimensioned at 1 MB and k=8 hash functions, cumulatively for query and insert, will take 12.4 $\mu$$s$ ($2\times(2.9+3.3)$ $\mu$$s$) on a Raspberry Pi 4 and 14.4 $\mu$$s$ on a Jetson Nano, on average.